OpenStudy (anonymous):

Let f(x) = cos^-1 x. Then the domain of f(x) is all real numbers and the range is [-1,1]. Is this true or false?

OpenStudy (anonymous):

It's false.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

\[ \cos : \mathbb{R}\mapsto [-1,1] \]\[ \cos^{-1} :[-1,1]\mapsto \left[\frac{-\pi}{2},\frac{\pi}{2}\right] \]

OpenStudy (anonymous):

what about if f(x) was tan^-1x then the domain(x) is all real numbers and the range was [0,pi], would in this case it be true?